You place a box weighing 218.5 N on an inclined plane that makes a 37.3◦ angle with the
horizontal.
Compute the component of the gravitational force acting down the inclined plane.
Answer in units of N.
Don't round answer.
1
Expert's answer
2019-10-15T06:19:54-0400

The component of the gravitational force acting down the inclined plane


"F=W\sin\theta=218.5\:\rm N\times\sin(37.3^{\circ})=132.408\:\rm N"